OpenStudy (anonymous):

working three days per week a clerk earns $255 per week.At the same rate of pay which expressionbelow represents what she would earn if she worked 5 days per week?(Let X represent full time earrnings.)

OpenStudy (abbles):

Right now, she is earning 85 dollars a day because 255 divided by 3 (the number of days she is working per week) is 85. Next, you would just multiply 85 by 5 to get 425. If she worked 5 days a week, she would earn $425 dollars a week.

OpenStudy (anonymous):

like the answers that are multiple choice are X=3/5(255) X=5/3(255) X=3(255) X=5(255) Not enough information is given